Table 1 shows the height and growth of two friends for a year.

Kamal = Height:145 cm Growth: 30%

Salman = Height: 140 cm Growth: 35%

By applying the concept of proportion, who is more higher after a year? Justify your choice.


1
Expert's answer
2021-08-20T09:03:12-0400

Kamal: "145\cdot(1+0.3)=188.5" cm.


Salmah: "140\cdot(1+0.35)=189" cm.


"189>188.5"

 Salmah is more higher after a year.